OpenStudy (prominer10000):

Which ordered pair is the solution to the system of equations? 2x-5y=-5 x+2y=11 A. (3, 4) B. (5, 3) C. (7, 2) D. (15, 7) @eliesaab

OpenStudy (bethb):

for the first equation i got -5 for two choices and i got 11 from two choices too

OpenStudy (eliesaab):

2x-5y=-5 x+2y=11 2x-5y=-5 -2x -4 y =-22 Add -9y =-27 y=3

OpenStudy (eliesaab):

You can now choose the answer

OpenStudy (prominer10000):

Is it B?

OpenStudy (eliesaab):

Yes

OpenStudy (prominer10000):

Ok thanks!